Circular No. 8490

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005am IN NGC 2811 R. Martin, Perth Observatory, reports the discovery of an apparent su-

pernova (mag 17.5) in the course of the Perth Automated Supernova Search on CCD images taken on Feb. 22.733 and 24.579 UT. The new object is located at α = 9^{h}16^{m}12^{s}.47, δ = −16^{o}18′16′′.0 (equinox 2000.0), which is 17′′.6 east and 31′′ north of the center of NGC 2811. Nothing was visible at this location on an image taken on Feb. 18.622 (limiting mag 18.5). Martin adds that the supernova had brightened between Feb. 22 and 24. H. Yamaoka, Kyushu University, writes that K. Itagaki (Teppo-cho, Yamagata, Japan) found SN 2005am at (red) mag 14.0 on a confirming unfiltered CCD image taken on Mar. 2.574 with a 0.60-m reflector, providing position end figures 13^{s}.07, 16′′.4 (adding that there is a foreground star ∼ 7′′ southwest of the new object at position end figures 12^{s}.88, 22′′.1).

SUPERNOVA 2005ah M. Salvo, J. Rich, and B. Schmidt, Australian National University

(ANU), report that a low-S/N spectrum (range 360–750 nm) of SN 2005ah (cf. IAUC 8486), obtained with the ANU 2.3-m telescope (+ Double-Beam Spectrograph) at Siding Spring Observatory on Feb. 18.48 UT, is consistent with that of a type-Ia supernova a few weeks after maximum light.

COMET C/2005 D1 (SOHO)

Continuation to IAUC 8488 (not a known group member):

Comet 2005 UT α_{2000} δ_{2000} Inst. F MPEC
C/2005 D1 Feb. 23.205 22^{h}22.^{m}2 −11^{o}52′ C2 HS 2005-E01

COMET C/1995 O1 (HALE-BOPP) S. Nakano, Sumoto, Japan, has acquired astrometry and photome-

try of this comet from unfiltered CCD images obtained by M. Tsumura (Wakayama, Japan) with a 0.20-m reflector and a 0.11-m refractor during visits to Waddi Farm in Western Australia, yielding the following total magnitudes for the comet (which showed a very diffuse coma of diameter ∼ 6′′): 2004 Mar. 21.63 UT, 17.8–17.9; 25.56, 17.8–18.0; 2005 Feb. 8.6, 18.3–18.4. There have been very few observations of this comet, now at r = 21 AU, in the last two years (see also IAUC 8479).

2005 March 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green